# Density matrix analysis of systems influenced by periodic Hamiltonians

### v1

@@ III Two-level system influenced by a time-dependent magnetic field
with $f_{0}\in\mathbb{R}^{+}$. The periodic square pulse $f(t)$ can be expressed as
$f(t)=f_{0}\Theta\left(\frac{T}{2}-t\right)-f_{0}\Theta\left(t-\frac{T}{2}\right)-f_{0}\delta_{t,\frac{T}{2}}$ (56)
with $\delta_{t,\frac{T}{2}}$ denoting Kronecker-delta. We consider the density matrix of the system as
@@
As before, from the first two equations, we obtain $\rho_{00}=1-\rho_{11}$. We again consider the state to be purely in the ground state such that $\rho_{00}(0)=1$ and $\rho_{01}(0)=\rho_{10}(0)=\rho_{11}(0)=0$. Multiplying eq.(s)(60,61) with $f(t)$ and summing them, we obtain a new equation of the form
@@
where in the second line of the above equation, we have made use of eq.(58). From eq.(56), we obtain that $\dot{f}(t)=-2f_{0}\delta\left(t-\frac{T}{2}\right)$. Substituting the expression of $\dot{f}(t)$ in the last line of eq.(62) and executing the time integral from 00 to $t$, we obtain
$\begin{split}\rho_{00}(t)=&\rho_{00}(0)-\frac{f(t)}{2}\left(\rho_{01}(t)+\rho_{10}(t)\right)\\ &-f_{0}\Theta(t-\frac{T}{2})\left(\rho_{01}(T/2)+\rho_{10}(T/2)\right)~{}.\end{split}$ (63)
@@
Subtracting eq.(61) from eq.(60) and multiplying both sides by $-if(t)E_{0}$, we obtain a new equation of the form
@@
The left hand side of the above equation can be expressed as
@@
$-if(t)E_{0}\left(\dot{\rho}_{01}(t)-\dot{\rho}_{10}(t)\right)=\ddot{\rho}_{00}(t)~{}.$ (69)
@@
The above equation is very elegant in a sense that for $t<\frac{T}{2}$ and $t>\frac{T}{2}$, eq.(70) reduces to the same structure given as
@@
where $\mathcal{E}_{0}\equiv E_{0}\sqrt{1+f_{0}^{2}}$. At $t=0$, $\rho_{00}(0)=1$ and $\dot{\rho}_{00}=0$. These initial conditions, help us to completely determine the anlytical form of $\rho_{00}(t)$ for $t<\frac{T}{2}$ as
@@
We now need to obtain the off-diagonal elements of the density matrix for which we need to make use of eq.(s)(58,66). For $t<\frac{T}{2}$ and using the solution of $\rho_{00}(t)$ from eq.(76), we can recast eq.(s)(58,66) as
@@
Again, using the condition $\rho_{01}(T/2)=\rho_{10}(T/2)=0$ we get back the same periodicity condition obtained earlier in eq.(75). One important thing to observe from eq.(s)(80,81) and eq.(s)(82,83) that the off-diagonal elements change there respective sign when the time crosses $t=\frac{T}{2}$. This is solely due to the form of the square pulse used in eq.(55) where $f(t)$ goes from $f_{0}$ for $t<T/2$ to $-f_{0}$ for $t\geq\frac{T}{2}$. The periodicity condition implies that at $t=\frac{T}{2},T,\frac{3T}{2},\cdots$ the density matrix represents a pure state $\rho=\begin{pmatrix}1&&0\\ 0&&0\end{pmatrix}$. In order to extract out the state of the system, we take a similar ansatz as before
@@
$\begin{split}\langle\phi_{t<T/2}|i\frac{\partial}{\partial}-\hat{H}|\phi_{t<T/2}\rangle&=0\\ \implies\dot{\theta}_{t<T/2}(t)&=0\\ \implies\theta_{t<T/2}(t)=\theta_{0}\end{split}$ (87)
@@ III.1.1 $l_{1}$-norm measure of coherence
In our model system $d=2$ and in our current analysis, we obtain the $l_{1}$-norm measure of coherence as
@@
We observe from Fig.(1) that in the first cycle at the starting time and at $t=T/2$, the measure of coherence is zero whereas it becomes maximum at $t=T/4$ when $f_{0}=0.1$. If $f_{0}$ is greater than unity then we observe a more interesting behaviour of the coherence measure as can be seen from Fig.(2). We observe from Fig.(2) that the measure of coherence increases then drops at $t=T/4$ and again increases and goes to zero at $t=\frac{T}{2}$. This same pattern follows for $\frac{T}{2}\leq t\leq T$. We also observe that the $l_{1}$-norm measure of coherence becomes unity for $f_{0}\geq 1$ as can be observed from Fig.(2). As can be seen from the expression in eq.(89), the coherence among the two states of the two-level system becomes zero provided $f_{0}=0$ indicating the time dependent part of the Hamiltonian is reponsible for generating this time-dependent coherence between the two states. This result is a bit counter-intuitive as from eq.(s)(85,86), we already find out that the density matrix is indeed separable for all value of $t$ and as a result there is no true mixing of states. To investigate this issue, we calculate the Frobenius norm measure of coherence.
@@ III.1.2 Frobenius-norm measure of coherence
where $\lambda_{1}=\lambda_{+}$ and $\lambda_{2}=\lambda_{-}$. Substituting the eigenvalues in eq.(91), we obtain the analytical form of Frobenius-norm measure of coherence as
@@
For both $t<T/2$ and $t>T/2$, $|\rho_{01}|^{2}$ has the same analytical structure as can be noted by observing the analytical forms of the off-diagonal elements of the density matrix in these two time sectors from eq.(s)(80,81,82,83). It is easy to check that $|\rho_{01}|^{2}-\rho_{00}\rho_{11}=0$ from the analytical form of the density matrix. Hence, the Frobenius-norm measure of coherence is
@@
The above result indeed is very interesting. The unity value of the Frobenius norm measure indicates that the system has geometrically the maximum distance from the maximally mixed state. Hence, the state is always coherent from a geometrical point of view. The primary observation lies in the fact that the $l_{1}$-norm measure of coherence is time-dependent whereas the Frobenius-norm measure of coherence is time-independent. This implies that the states are fully coherenet from a geometric point of view which also coincides with the outcome that the density matrix is separable for all values of $t$. Here, $l_{1}$-norm coherence measure indeed does not capture the full picture as the state is completely separable at $t=0,T/2,\frac{3T}{2}\cdots$, indicating that it is very away from the maximally mixed state which is the two dimensional identity matrix with appropriate numerical pre-factor. Hence, for such a scenario Frobenius-norm measure of coherence is more reliable than the $l_{1}$-norm measure of coherence.
@@ IV Conclusion
In this work, we have considered interaction of two-level systems with time-dependent Hamiltonians which are periodic in time. We resort to a density matrix approach, where we aim to solve the Liouville equations of motion to obtain the analytical form of the density matrix. At first, we solve exactly the standard Rabi-oscilation model by solving the four independent differential equations corresponding to the four components of the density matrix. Making use of appropriate initial conditions and the Floquet theorem, we then obtain the state of the system. Next, we have made use of the Lewis-Riesenfeld invariant approach to solve the similar model. The interesting observation lies in the fact the Lewis invariant operator has the same analytical structure of the density matrix solved in the initial analysis. This gives us a very robust connection between the density marix as well as the Lewis invariant approach. We therefore conclude that the density matrix of a system is the same as the Lewis invariant operator for a system upto some constant factor and we also find out that the Lewis phase of the system for a Hamiltonian with periodicity is some constant multiplied by time. Finally, using this density matrix approach, we solve a two-level system influenced by a time-independent constant magnetic magnetic field in the $z$-direction and a time-dependent magnetic field in the $x$-direction. For a periodic square pulse, we observe a very interesting solution of the density matrix where the diagonal elements remain same over time and the off-diagonal elements pick up a negative sign when $t$ crosses $T/2$. From the value of the elements of the density matrix at $t=T/2$, we obtain a periodicity condition. We find out that the density matrix is separable for all times for this two-level system influenced by a periodic magnetic field and the state has a different structure for $t<T/2$ and $t>T/2$. We also find out that for $t>T/2$, the coefficient corresponding to state $|1\rangle$ changes it sign. We have then plotted the $l_{1}$-norm measure of coherence which depicts an explicit time dependence and the coherence becomes maximum ($C_{l_{1}}(\rho)=1$) for $f_{0}\geq 1$. On the other hand, if we compute the Frobenius-norm measure of coherence, we find out that the measure of coherence is unity and independent of time indicating a fully coherent state throughout which is more reliable as we have found that the density matrix is separable for all values of $t$. Hence, it is interesting to find out that for such a scenario the Frobenius-norm measure of coherence which is a geometric interpretation of the distance of a state from the maximally entangled state depicts the true nature of coherence implementing that the $l_{1}$-norm measure of coherence may not be reliable always.

### v2

@@ III Two-level system influenced by a time-dependent magnetic field
with $f_{0}\in\mathbb{R}^{+}$. The periodic square pulse $f(t)$ in the above equation can be written using the Heaviside theta function as
$\begin{split}f(t)=&f_{0}\Theta\left(\frac{T}{2}-t\right)-f_{0}\Theta\left(t-\frac{T}{2}\right)\\ &-4f_{0}\Theta\left(\frac{T}{2}-t\right)\Theta\left(t-\frac{T}{2}\right)\end{split}$ (56)
where the Heaviside theta function $\Theta\left(t-\frac{T}{2}\right)$ is defined as
$\Theta\left(t-\frac{T}{2}\right)=\left\{\begin{matrix}1&&t>\frac{T}{2}\\ 0&&t<\frac{T}{2}\\ \frac{1}{2}&&t=\frac{T}{2}\end{matrix}\right.$~. (57)
We consider the density matrix of the system as
@@
As before, from the first two equations, we obtain $\rho_{00}=1-\rho_{11}$. We again consider the state to be purely in the ground state such that $\rho_{00}(0)=1$ and $\rho_{01}(0)=\rho_{10}(0)=\rho_{11}(0)=0$. Multiplying eq.(s)(61,62) with $f(t)$ and summing them, we obtain a new equation of the form
@@
where in the second line of the above equation, we have made use of eq.(59). From eq.(56), we can calculate the analytical form of the time derivative of the function $f(t)$ as
$\begin{split}\dot{f}(t)=&-2f_{0}\delta\left(\frac{T}{2}-t\right)-4f_{0}\biggr[\Theta\left(\frac{T}{2}-t\right)\delta\left(t-\frac{T}{2}\right)\\ &-\delta\left(t-\frac{T}{2}\right)\Theta\left(t-\frac{T}{2}\right)\biggr]\\ =&-2f_{0}\delta\left(t-\frac{T}{2}\right)-2f_{0}\left[\delta\left(t-\frac{T}{2}\right)-\delta\left(t-\frac{T}{2}\right)\right]\\ =&-2f_{0}\delta\left(t-\frac{T}{2}\right)~.\end{split}$ (64)
Substituting the expression of $\dot{f}(t)$ in the last line of eq.(63) and executing the time integral from 0 to $t$, we obtain
$\begin{split}\rho_{00}(t)=&\rho_{00}(0)-\frac{f(t)}{2}\left(\rho_{01}(t)+\rho_{10}(t)\right)\\ &-f_{0}\Theta\left(t-\frac{T}{2}\right)\left(\rho_{01}(T/2)+\rho_{10}(T/2)\right)~.\end{split}$ (65)
@@
Subtracting eq.(62) from eq.(61) and multiplying both sides by $if(t)E_{0}$, we obtain a new equation of the form
@@
The left-hand side of the above equation can be expressed as
@@
$if(t)E_{0}\left(\dot{\rho}_{01}(t)-\dot{\rho}_{10}(t)\right)=\ddot{\rho}_{00}(t)~.$ (71)
@@
The above equation is very elegant in the sense that for $t<\frac{T}{2}$ and $t>\frac{T}{2}$, eq.(72) reduces to the same structure given as
@@
where $\mathcal{E}_{0}\equiv E_{0}\sqrt{1+f_{0}^{2}}$. At $t=0$, $\rho_{00}(0)=1$ and $\dot{\rho}_{00}=0$. These initial conditions help us to completely determine the analytical form of $\rho_{00}(t)$ for $t<\frac{T}{2}$ as
@@
We now need to obtain the off-diagonal elements of the density matrix for which we need to make use of eq.(s)(59,68). For $t<\frac{T}{2}$ and using the solution of $\rho_{00}(t)$ from eq.(78), we can recast eq.(s)(59,68) as
@@
Again, using the condition $\rho_{01}(T/2)=\rho_{10}(T/2)=0$ we get back the same periodicity condition obtained earlier in eq.(77). One important thing to observe from eq.(s)(82,83) and eq.(s)(84,85) is that the off-diagonal elements change their respective sign when the time crosses $t=\frac{T}{2}$. This is solely due to the form of the square pulse used in eq.(55) where $f(t)$ goes from $f_{0}$ for $t<T/2$ to $f_{0}$ for $t\geq\frac{T}{2}$. The periodicity condition implies that at $t=\frac{T}{2},T,\frac{3T}{2},\cdots$ the density matrix represents a pure state $\rho=\begin{pmatrix}1&&0\\ 0&&0\end{pmatrix}$. In order to extract the state of the system, we take a similar ansatz as before
@@
$\begin{split}\langle\phi_{t<T/2}|i\frac{\partial}{\partial t}-\hat{H}|\phi_{t<T/2}\rangle&=0\\ \implies\dot{\theta}_{t<T/2}(t)&=0\\ \implies\theta_{t<T/2}(t)&=\theta_{0}\end{split}$ (89)
@@ III.1.1 $l_{1}$-norm measure of coherence
In our model system, $d=2$ and in our current analysis, we obtain the $l_{1}$-norm measure of coherence as
@@
We observe from Fig.(1) that in the first cycle at the starting time and at $t=T/2$, the measure of coherence is zero, whereas it becomes maximum at $t=T/4$ when $f_{0}=0.1$. If $f_{0}$ is greater than unity, then we observe a more interesting behaviour of the coherence measure as can be seen from Fig.(2). We observe from Fig.(2) that the measure of coherence increases, then drops at $t=T/4$ and again increases and goes to zero at $t=\frac{T}{2}$. This same pattern follows for $\frac{T}{2}\leq t\leq T$. We also observe that the $l_{1}$-norm measure of coherence becomes unity for $f_{0}\geq 1$ as can be observed from Fig.(2). As can be seen from the expression in eq.(91), the coherence among the two states of the two-level system becomes zero provided $f_{0}=0$, indicating the time-dependent part of the Hamiltonian is responsible for generating this time-dependent coherence between the two states. This result is a bit counter-intuitive as from eq.(s)(87,88), we already find out that the density matrix is indeed separable for all values of $t$ and as a result there is no true mixing of states. To investigate this issue, we calculate the Frobenius norm measure of coherence.
@@ III.1.2 Frobenius-norm measure of coherence
where $\lambda_{1}=\lambda_{+}$ and $\lambda_{2}=\lambda_{-}$. Substituting the eigenvalues in eq.(93), we obtain the analytical form of the Frobenius-norm measure of coherence as
@@
For both $t<T/2$ and $t>T/2$, $|\rho_{01}|^{2}$ has the same analytical structure as can be noted by observing the analytical forms of the off-diagonal elements of the density matrix in these two time sectors from eq.(s)(82,83,84,85). It is easy to check that $|\rho_{01}|^{2}-\rho_{00}\rho_{11}=0$ from the analytical form of the density matrix. Hence, the Frobenius-norm measure of coherence is
@@
The above result is indeed very interesting. The unity value of the Frobenius norm measure indicates that the system has geometrically the maximum distance from the maximally mixed state. Hence, the state is always coherent from a geometrical point of view. The primary observation lies in the fact that the $l_{1}$-norm measure of coherence is time-dependent, whereas the Frobenius-norm measure of coherence is time-independent. This implies that the states are fully coherent from a geometric point of view, which also coincides with the outcome that the density matrix is separable for all values of $t$. Here, $l_{1}$-norm coherence measure indeed does not capture the full picture as the state is completely separable at $t=0,T/2,\frac{3T}{2}\cdots$, indicating that it is very far from the maximally mixed state, which is the two-dimensional identity matrix with appropriate numerical pre-factor. Hence, for such a scenario, the Frobenius-norm measure of coherence is more reliable than the $l_{1}$-norm measure of coherence.
@@ IV Conclusion
In this work, we have considered the interaction of two-level systems with time-dependent Hamiltonians which are periodic in time. We resorted to a density matrix approach, where our primary aim was to solve the von Neumann equations of motion to obtain the analytical form of the density matrix. At first, we solved exactly the standard Rabi-oscillation model by solving the four independent differential equations corresponding to the four components of the density matrix. Making use of appropriate initial conditions and the Floquet theorem, we then obtain the state of the system. Next, we have made use of the Lewis-Riesenfeld invariant approach to solve the same model. The interesting observation lies in the fact that the Lewis invariant operator has the same analytical structure as that of the density matrix solved in the initial analysis. This gives us a very robust connection between the density matrix as well as the Lewis invariant approach. We therefore conclude that the density matrix of a system is the same as the Lewis invariant operator for a system up to some constant factor, and we also find out that the Lewis phase of the system for a Hamiltonian with periodicity is some constant multiplied by time. Finally, using this density matrix approach, we went on to solve a two-level system influenced by a time-independent constant magnetic field in the $z$-direction and a time-dependent magnetic field in the $x$-direction. For a periodic square pulse, we observe a very interesting solution of the density matrix where the diagonal elements remain the same over time and the off-diagonal elements pick up a negative sign when $t$ crosses $T/2$. From the value of the elements of the density matrix at $t=T/2$, we obtain a periodicity condition. We find out that the density matrix is separable for all times for this two-level system influenced by a periodic magnetic field, and the state has a different structure for $t<T/2$ and $t>T/2$. We also find out that for $t>T/2$, the coefficient corresponding to state $|1\rangle$ changes its sign. We have then plotted the $l_{1}$-norm measure of coherence, which depicts an explicit time dependence, and the coherence becomes maximum ($C_{l_{1}}(\rho)=1$) for $f_{0}\geq 1$. On the other hand, if we compute the Frobenius-norm measure of coherence, we find that the measure of coherence is unity and independent of time, indicating a fully coherent state throughout, which is more reliable, as we have found that the density matrix is separable for all values of $t$. Hence, it is interesting to find out that for such a scenario, the Frobenius-norm measure of coherence, which is a geometric interpretation of the distance of a state from the maximally entangled state, depicts the true nature of coherence, implying that the $l_{1}$-norm measure of coherence may not be reliable always.
Acknowledgement
We thank the anonymous referee for the useful comments that have helped to improve our paper.
Data availability statement
This manuscript has no associated data, and no new data were generated during the completion of the manuscript.